a particle moves with a displacement function S = 20t - 1/2t^2. calculate the time taken for the particle to come to rest.
1
Expert's answer
2019-06-05T09:49:09-0400

velosiy is the desplacement derivation and equals to

"v = 20 - 0.5(-2)t^{-3} = 20 + t^{-3}"

so the particle will never come to rest, only at time aspirant to infinity...